# Instability and breaking of internal waves in a horizontal shear layer

### v1

@@ 7 Discussion and conclusions
Throughout this study, we have neglected the influence of planetary rotation that will undoubtedly be important when the time scales of either the shear layer or the internal wave are sufficiently long, that is, when $\Omega\sim O(f)$ or $\partial U/\partial y\sim O(f)$. The latter condition was studied for $N\sim f$ by staquet2002transport, who suggested that in the case of wave trapping, inertial instability may act as an additional source of turbulence energy. The former condition may be particularly relevant for scenarios where wave advection of momentum dominates energy perturbation growth, which is favoured for small horizontal wavenumbers $k_{y},k_{x}\ll 1$ and hence small $\Omega/N$. In this case, the dynamics in the presence of rotation may be considerably more complicated, as is suggested by the recent results of hilditch2025. The effect of changing the Reynolds number $Re_{w}$ will undoubtedly be important during wave breaking and the turbulence that follows, especially at the modest values considered here. In particular, the extents to which the critical layer dynamics (e.g. wang2020nonlinearii), secondary instabilities (e.g. Mashayeka13) and the dissipation and mixing rates shown in figure 16 are affected by changes in $Re_{w}$ remain open problems. A further parameter that continues to present a considerable computational challenge in direct numerical simulations is the Prandtl number $Pr$ that we have assumed to be unity, but in reality is expected to be $O(10)$ in thermally-stratified ocean environments.
@@
Finally, to what extent the variability in wave breaking dynamics explored here extends to a more general class of background flows–which might include cyclostrophically balanced flows or, for non-zero $f$, rotationally balanced flows (see, e.g. whitt2013near; edwards2005focusing)–remains an open question. Scenarios such as these, in addition to the problem studied here, may help to provide a crucial link between wave breaking and the asymptotic regime of anistropic stratified turbulence that has been proposed as a universal description for the dynamics of turbulent patches in the ocean interior (dasaro2022internal; kunze2019unified; riley2008stratified). As dasaro2022internal points out, the initial condition for the classical problem of a (horizontally infinite) wave packet incident on a vertical shear flow lacks potential vorticity, thought to be a key ingredient in the energy cascade of strongly stratified turbulent flows (lindborg2006energy). The inclusion of a horizontally varying mean flow provides a possible source of potential vorticity at large scales. It would be worthwhile to investigate whether the turbulence produced in the wave breaking scenarios studied here can access this strongly-stratified regime in the case when the mean flow is linearly stable (the unstable case is discussed in detail in lewin2024evidence). It is also worth recognising that, in general (and as was the case in this study), horizontally bounded wave packets carry with them their own source of potential vorticity in the form of a Lagrangian vorticity dipole (buhler2005wave; bretherton1969mean). This may, as suggested by dasaro2022internal and by the focussing of the mean flow shown in figure 12, play an important role in the resulting energy cascade to small scales during breaking.
@@
[Funding]MMPC acknowledges the Natural Sciences and Engineering Research Council of Canada (NSERC) through RGPIN-2024-06184. SFL and AKK were supported by the Ho-Shang and Mei-Li Lee Faculty Fellowship in Mechanical Engineering at UC Berkeley. AB acknowledges support from a National Science Foundation Graduate Research Fellowship (Grant Number DGE-1656518). This research used the Savio computational cluster resource provided by the Berkeley Research Computing program at the University of California, Berkeley (supported by the UC Berkeley Chancellor, Vice Chancellor for Research, and Chief Information Officer).
@@
[Author ORCIDs] S. Lewin, https://orcid.org/0000-0002-2602-4751; A. Kaminski, https://orcid.org/0000-0002-4838-2453; M. Couchman, https://orcid.org/0000-0002-4667-6829
Appendix A Inclined coordinates
@@
A schematic of the inclined coordinate system introduced in section §3.2 is shown in figure 18. Under rotation by an angle $\alpha=-\arctan(k_{x0}/k_{z0})$ in the xxzz plane, the velocities in the $\tilde{x}$ and $\tilde{z}$ directions are $(\hat{u},\hat{w})=(u\cos\alpha+w\sin\alpha,-u\sin\alpha+w\cos\alpha$). The equations of motion are
@@
Appendix B Wave-induced mean flow
As originally demonstrated by bretherton1969mean, the stress exerted by an internal wave packet on the fluid through which it propagates induces a mean flow at second order in the wave amplitude. The nature of this mean flow depends on the geometry of the wave packet, in particular the aspect ratio of its envelope. van2018wave provide a general expression for the wave-induced mean flow in a wave packet with arbitrary amplitude envelope $A(x,y,z,t)$, writing the result as the sum of the horizontal Bretherton flow forced by vertical vorticity and a long wave flow forced by horizontal vorticity. Because the wave packets we consider here are unbounded in the vertical, the long wave flow is zero at leading order. Extending the analysis of van2018wave to a 3D wave packet, horizontal divergence of the momentum flux drives a mean flow $\mathbf{u}_{DF}$ given by
@@ Appendix B Wave-induced mean flow
taking the form of the pseudomomentum per unit mass in the $x$-direction. This so-called divergent-flux induced flow must induce a secondary response flow to ensure incompressibility; the sum of these two flows gives the Bretherton flow $\mathbf{u}_{BF}$. As shown in van2018wave, $w_{BF}=0$ and we must have additionally that $(\nabla\times\mathbf{u}_{DF})\cdot\hat{\mathbf{k}}=(\nabla\times\mathbf{u}_{BF})\cdot\hat{\mathbf{k}}$. In combination with the incompressibility condition $\nabla\cdot\mathbf{u}_{BF}=0$, it follows that, in Fourier space with coordinates $(\kappa,\lambda,\mu)$,

### v2

@@ 7 Discussion and conclusions
Throughout this study, we have neglected the influence of planetary rotation that will undoubtedly be important when the time scales of either the shear layer or the internal wave are sufficiently long, that is, when $\Omega\sim O(f)$ or $\partial U/\partial y\sim O(f)$. The latter condition was studied for $N\sim f$ by Staquet and Huerre (2002), who suggested that in the case of wave trapping, inertial instability may act as an additional source of turbulence energy. The former condition may be particularly relevant for scenarios where wave advection of momentum dominates energy perturbation growth, which is favoured for small horizontal wavenumbers $k_{y},k_{x}\ll 1$ and hence small $\Omega/N$. In this case, the dynamics in the presence of rotation may be considerably more complicated, as is suggested by the recent results of Hilditch, Taylor, and Thomas (2025). The effect of changing the Reynolds number $Re_{w}$ will undoubtedly be important during wave breaking and the turbulence that follows, especially at the modest values considered here. In particular, the extents to which the critical layer dynamics (e.g. Wang and Balmforth, 2021), secondary instabilities (e.g. Mashayek and Peltier, 2013) and the dissipation and mixing rates shown in figure 16 are affected by changes in $Re_{w}$ remain open problems. A further parameter that continues to present a considerable computational challenge in direct numerical simulations is the Prandtl number $Pr$ that we have assumed to be unity, but in reality is expected to be $O(10)$ in thermally-stratified ocean environments.
@@
Finally, to what extent the variability in wave breaking dynamics explored here extends to a more general class of background flows–which might include cyclostrophically balanced flows or, for non-zero $f$, rotationally balanced flows (see, e.g. Whitt and Thomas, 2013; Edwards and Staquet, 2005)–remains an open question. Scenarios such as these, in addition to the problem studied here, may help to provide a crucial link between wave breaking and the asymptotic regime of anistropic stratified turbulence that has been proposed as a universal description for the dynamics of turbulent patches in the ocean interior (D’Asaro, 2022; Kunze, 2019; Riley and Lindborg, 2008). As D’Asaro (2022) points out, the initial condition for the classical problem of a (horizontally infinite) wave packet incident on a vertical shear flow lacks potential vorticity, thought to be a key ingredient in the energy cascade of strongly stratified turbulent flows (Lindborg, 2006). The inclusion of a horizontally varying mean flow provides a possible source of potential vorticity at large scales. It would be worthwhile to investigate whether the turbulence produced in the wave breaking scenarios studied here can access this strongly-stratified regime in the case when the mean flow is linearly stable (the unstable case is discussed in detail in Lewin and Caulfield (2024)). It is also worth recognising that, in general (and as was the case in this study), horizontally bounded wave packets carry with them their own source of potential vorticity in the form of a Lagrangian vorticity dipole (Bühler and McIntyre, 2005; Bretherton, 1969). This may, as suggested by D’Asaro (2022) and by the focussing of the mean flow shown in figure 12, play an important role in the resulting energy cascade to small scales during breaking.
@@
[Funding]MMPC acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC) through RGPIN-2024-06184. SFL and AKK were supported by the Ho-Shang and Mei-Li Lee Faculty Fellowship in Mechanical Engineering at UC Berkeley. AB acknowledges support from a National Science Foundation Graduate Research Fellowship (Grant Number DGE-1656518). This research used the Savio computational cluster resource provided by the Berkeley Research Computing program at the University of California, Berkeley (supported by the UC Berkeley Chancellor, Vice Chancellor for Research, and Chief Information Officer).
@@
[Author ORCIDs] S. Lewin, https://orcid.org/0000-0002-2602-4751; A. Kaminski, https://orcid.org/0000-0002-4838-2453; A. Balakrishna, https://orcid.org/0009-0006-2696-3130; M. Couchman, https://orcid.org/0000-0002-4667-6829
Appendix A Minimum gradient Richardson number
After some algebra, equation (22) for the gradient Richardson number can be written as
$Ri_{g}=\frac{R^{2}}{F^{2}s^{2}}\frac{1+s\cos\phi}{\left[\left(R\cos\phi+(k_{x}/k_{y})\sin\phi\right)^{2}+\sin^{2}\phi\right]\sin^{2}(\beta+\gamma)},$ (39)
where $\tan\gamma=\sin\phi/(R\cos\phi+k_{x}\sin\phi/k_{y})$. Thus
$\min_{\phi,\beta}(Ri_{g})=\frac{R^{2}}{F^{2}s^{2}}\min_{\phi}\left(\frac{1+s\cos\phi}{\left(R\cos\phi+(k_{x}/k_{y})\sin\phi\right)^{2}+\sin^{2}\phi}\right).$ (40)
In the following we will assume $s<1$ so that $Ri_{g}>0$. Consider first the case $k_{x}=0$. For $R\geq 1$, the denominator and numerator of the function in brackets are respectively maximised and minimised at the same value of $\phi=\pi$, making this a global minimiser. For $R<1$, let $c=\cos\phi$, then stationary points of the function in brackets above are given by the roots of a quadratic equation:
$c_{min}=\frac{1}{s}\left(-1\pm\sqrt{1-s^{2}/(1-R^{2})}\right).$ (41)
There are two cases to consider. When $1-R^{2}<s^{2}$, the discriminant in the expression above is negative and there are no stationary points in the range 1<c<1-1<c<1. The global minimum thus occurs at $c=-1\iff\phi=\pi$. On the other hand, when $1-R^{2}>s^{2}$ there is a single global minimum in the range 1<c<1-1<c<1 at $c=c_{min}$. Substituting in we obtain:
[…]
Note in particular that $R^{2}/F^{2}>1=$ so that, when $R^{2}\leq 1-s^{2}$, $\min(Ri_{g})>1/2$ for $s<1$.
When $k_{x}\neq 0$, the denominator of (40) is not maximised at $\phi=\pi/2$ for any value of $R$. However, the value at $\phi=\pi/2$ will be similar to the maximum value when $Rk_{y}/k_{x}$ is large. Because $Rk_{y}/k_{x}=Fk_{h}|\mathbf{k}|/(k_{x}k_{z})>F$, we can argue heuristically that this may be anticipated near locations of wave instability. Where dynamics are dominated by wave steepening near trapping levels, $k_{h}|\mathbf{k}|/(k_{x}k_{z})\gg 1$. On the other hand, where vertical shear produced by wave advection of momentum dominates, $F$ is necessarily large. Figure 18 demonstrates that, indeed, $(1-s)/F^{2}s^{2}$ becomes an increasingly good approximation for $\min(Ri_{g})$ as the latter decreases towards zero. Note that, in particular, the point at which $\min(Ri_{g})$ becomes smaller than $1/4$ is very well estimated using $(1-s)/F^{2}s^{2}$ in all cases.
Figure 18: Evolution of the minimum gradient Richardson number along rays for the parameter values considered in figure 2, computed by evaluating (40) numerically. Dotted curves show the estimate $\min(Ri_{g})=(1-s)/F^{2}s^{2}$. The heuristic local stability threshold $\min(Ri_{g})=1/4$ is also indicated.
Appendix B Inclined coordinates
@@
A schematic of the inclined coordinate system introduced in section §3.2 is shown in figure 19. Under rotation by an angle $\alpha=-\arctan(k_{x0}/k_{z0})$ in the $x$-$z$ plane, the velocities in the $\tilde{x}$ and $\tilde{z}$ directions are $(\hat{u},\hat{w})=(u\cos\alpha+w\sin\alpha,-u\sin\alpha+w\cos\alpha$). The equations of motion are
@@
Appendix C Wave-induced mean flow
As originally demonstrated by Bretherton (1969), the stress exerted by an internal wave packet on the fluid through which it propagates induces a mean flow at second order in the wave amplitude. The nature of this mean flow depends on the geometry of the wave packet, in particular the aspect ratio of its envelope. Van den Bremer and Sutherland (2018) provide a general expression for the wave-induced mean flow in a wave packet with arbitrary amplitude envelope $A(x,y,z,t)$, writing the result as the sum of the horizontal Bretherton flow forced by vertical vorticity and a long wave flow forced by horizontal vorticity. Because the wave packets we consider here are unbounded in the vertical, the long wave flow is zero at leading order. Extending the analysis of Van den Bremer and Sutherland (2018) to a 3D wave packet, horizontal divergence of the momentum flux drives a mean flow $\mathbf{u}_{DF}$ given by
@@ Appendix C Wave-induced mean flow
taking the form of the pseudomomentum per unit mass in the $x$-direction. This so-called divergent-flux induced flow must induce a secondary response flow to ensure incompressibility; the sum of these two flows gives the Bretherton flow $\mathbf{u}_{BF}$. As shown in Van den Bremer and Sutherland (2018), $w_{BF}=0$ and we must have additionally that $(\nabla\times\mathbf{u}_{DF})\cdot\hat{\mathbf{k}}=(\nabla\times\mathbf{u}_{BF})\cdot\hat{\mathbf{k}}$. In combination with the incompressibility condition $\nabla\cdot\mathbf{u}_{BF}=0$, it follows that, in Fourier space with coordinates $(\kappa,\lambda,\mu)$,